# Graph-based Confidence Calibration for Large Language Models

### v1

@@ 4.3 Out of Domain evaluation
Experiment setup: We evaluate the confidence calibration of different approaches under out-of-domain settings. We have two experiment configurations: out-of-domain dataset OODD, and out-of-domain LLMs (OODL). For OODD, we train the confidence calibration model on TriviaQA from Llama3 responses and test it on CoQA Llama3 answers. For OODL, we use the same training data from Llama3 but test the Vicuna model’s responses on the TriviaQA and CoQA datasets. We compare our model with the Apricot and GraphSpectral (with Platt scaling) methods.
Results and Analysis: The OOD experiments results revealed that our model maintained a high level of performance across tested domains Table. 3. Specifically, the model demonstrated consistent calibration, as evidenced by low ECE values and strong discriminative ability, reflected in high AUROC scores on in-domain and OOD datasets. For example, while the model achieved an ECE of 0.025 and an AUROC of 0.84 on TriviaQA (in-domain), it maintained an ECE of 0.077 and an AUROC of 0.77 on CoQA. Furthermore, the Brier scores across domains remained within acceptable ranges, demonstrating reliable probabilistic predictions even when faced with unfamiliar data distributions. The relatively small increase in ECE and a slight decrease in AUROC for OOD datasets suggest that while there is some degradation in performance, the model retains substantial robustness and accuracy. This is primarily because similarity graph patterns are highly invariant to the data distribution. Specifically, our model employs the consistency graph and the clustering feature that does not alter with data distribution shifts, enabling it to maintain stable performance across different datasets. In contrast, Apricot typically relies on specific dataset features, which leads to poor performance in OOD scenarios. Furthermore, calibration methods like the Platt scaling can improve the confidence calibration in-domain, but their calibration effectiveness remains limited under domain shift scenarios. This is because this calibration technique mainly adjusts the output probabilities but does not fundamentally address the biases introduced by feature representation changes across distributions. Table. 3 shows the OOD performance of the baseline methods.
Table 3: OOD evaluation of the graph spectral(w Platt scaling) trained on the TriviaQA from Llama3 responses and tested out-of-domain datasets
@@
llama3 CoQA GraphSpectral(w Platt scaling) 0.17 0.72 0.095
@@
Vicuna TriviaQA GraphSpectral(w Platt scaling) 0.24 0.53 0.07
@@
Vicuna CoQA GraphSpectral(w Platt scaling) 0.35 0.55 0.26
@@ 4.4 Sensitivity Analysis
Number of training samples We conducted experiments to examine the relationship between performance and the amount of training data. Specifically, we tested our model performance on the Llama3 TriviaQA dataset and varied the training size from 100 to 4000. The results are displayed in Table  4. We observed that the model’s performance does not drop significantly with the reduced training data. These experimental results indicate that the model performs well with limited data availability, demonstrating its applicability in real-world scenarios where only smaller datasets are available. We also tested the baseline performance, the results are shown in Appendix A.5.
@@
# of Training Samples ECE AUROC Brier
100 0.095 0.770 0.201
300 0.062 0.786 0.187
500 0.049 0.792 0.181
1000 0.037 0.799 0.177
4000 0.020 0.820 0.172
Hyperparameter sensitivity We conduct the sensitivity analysis of our model’s calibration error performance concerning two key configurations: the number of sampled answers used to construct the graph and the number of Graph Convolutional Network (GCN) layers in the GNN model. The results are displayed in Fig. 3. The experiments are conducted using the Llama3 model on the TriviaQA dataset. For Fig. 3 (a) experiments, we varied the number of sampled answers from 10 to 50 while keeping other configurations and hyperparameters fixed, as described in the experimental setup. We observe that increasing the number of sampled answers slightly improves performance, which then stabilizes. In Fig. 3(b), the sensitivity to the number of GCN layers indicates that our model remains stable with 1 to 4 layers, with the best performance observed at three layers.
@@ 5 Conclusion and Future Work
In summary, in this work, we proposed one effective strategy of confidence calibration by combining the LLM’s self-consistency with labeled data and training an auxiliary GNN model to estimate the correctness of its responses to questions. Experiments demonstrate that the proposed approach improves confidence calibration significantly across several datasets compared to baseline methods. We acknowledge that one limitation of our approach is the computation time of sampling multiple answers from LLMs. However, this cost can be effectively alleviated by parallel sampling. In the Appendix A.2, we provide the details of the sampling time. In future work, we aim to extend the framework to incorporate the data uncertainty coming from ambiguous questions and also explore the multi-step confidence calibration in the chain-of-thought framework. Our approach relies on a consistency graph to calibrate confidence.
Large Language Models (LLMs) are powerful tools that deliver valuable information across a wide range of applications. However, the misinformation they may generate can lead to significant societal issues. Our calibration model evaluates the accuracy of an LLM’s responses, thereby enhancing their reliability. When integrated into real-world systems, our model would enable LLMs to abstain from answering uncertain queries or allow users to decide the level of trust they place in the responses. We anticipate that our research will positively influence the responsible deployment and use of LLMs in society.
Appendix A Appendix
A.1 Hyperparameters and Model configurations
@@
Our model used three GCN layers; typically, the embedding dimension was $1024256,512,and$ 1024 for three GCN layers. For the training process, we used the binary cross-entropy loss with a decaying learning rate that reduced the learning rate by $0.9$ if the validation loss did not improve $10$ epochs (with an initial learning rate of $410^{-}$ 4 and a minimum learning rate of $10^{-7}$). The optimizer was Adam with $\beta_{1}=0.9$ and $\beta_{2}=0.98$. The batch size was 32. For the rephrased prompts, we set $k=3$, $n=30$, so for each rephrased question, we sampled ten answers. While calculating the ECE, we divide the confidence into $B=10$ bins.
@@ A.2 Computational cost
A.3 Additional Cases
To better understand our method intuitively, we have collected a few examples to show the difference between our algorithm and APRICOT.
To summarize our observation here:
@@ A.3 Additional Cases
Self-evaluation: $80\%$ (with the further question ”Please provide your confidence in the answer only in percent (0-100 %):”)
@@ A.5 Additional results
Figure 5: Reliability diagrams for different methods using ten bins each for TriviaQA from Llama3 model responses. The color and the percentage number within each bar indicates the ratio of responses contained in each bin. Larger values are represented by colors closer to purple.
Figure 6: Reliability diagrams for different methods using ten bins each for CoQA from Llama3 model responses. The color and the percentage number within each bar indicates the ratio of responses contained in each bin. Larger values are represented by colors closer to purple.
@@
A.6 Prompting strategy
Here, we showed the prompts we used to generate the rephrasing questions.
You are a helpful assistant. I have a question that I would like to see it rephrased in multiple ways. Please take the original question and generate several rephrased versions while maintaining the same meaning. The question can only have one direct answer. Here is the original question: …. Please provide four distinct rephrases of the question.

### v2

@@ 4.3 Confidence calibration in the OOD setting
Experiment setup: We evaluate the confidence calibration of different approaches under out-of-domain settings. We have two experiment configurations: out-of-domain dataset OODD, and out-of-domain LLMs (OODL). For OODD, we train the confidence calibration model on TriviaQA from Llama3 responses and test it on CoQA Llama3 and TruthfulQA Llama3 answers. For OODL, we use the same training data from Llama3 but test the Vicuna model’s responses on the TriviaQA and CoQA datasets. We compare our model with the Apricot and GraphSpectral (with Platt scaling) methods.
Results and Analysis: Table. 3 shows the OOD performance of the baseline methods. The OOD experiment results revealed that our model maintained a high level of performance across tested domains. Specifically, the model demonstrated consistent calibration, as evidenced by low ECE values and strong discriminative ability, reflected in high AUROC scores on in-domain and OOD datasets. For example, while the model achieved an ECE of 0.03 and an AUROC of 0.86 on TriviaQA (in-domain), it maintained an ECE of 0.077 and an AUROC of 0.77 on CoQA. Furthermore, the Brier scores across domains remained within acceptable ranges, demonstrating reliable probabilistic predictions even when faced with unfamiliar data distributions. The relatively small increase in ECE and a slight decrease in AUROC for OOD datasets suggest that while there is some degradation in performance, the model retains substantial robustness and accuracy. This is primarily because similarity graph patterns are highly invariant to the data distribution. Specifically, our model employs the consistency graph and the clustering feature that does not alter with data distribution shifts, enabling it to maintain stable performance across different datasets.
Table 3: Evaluation in the OOD setting. Models are trained on the TriviaQA from Llama3 responses and tested on out-of-domain datasets
@@
Llama3 CoQA GraphSpectral(w platt) 0.17 0.72 0.095
@@
Llama3 TruthfulQA GraphSpectral(w platt) 0.32 0.63 0.324
Apricot 0.25 0.54 0.197
Ours 0.23 0.66 0.16
Vicuna TriviaQA GraphSpectral(w platt) 0.24 0.53 0.07
@@
Vicuna CoQA GraphSpectral(w platt) 0.35 0.55 0.26
@@
In contrast, Apricot typically relies on specific dataset features, which leads to poor performance in OOD scenarios. Furthermore, calibration methods like the Platt scaling can improve the confidence calibration in-domain, but their calibration effectiveness remains limited under domain shift scenarios. This is because this calibration technique mainly adjusts the output probabilities but does not fundamentally address the biases introduced by feature representation changes across distributions.
4.4 Checking the quality automatic labels
In our previous experiment, we used an automated method to assign the correctness labels to responses by comparing them to true answers with ROUGE-L scores. One question is whether these labels are reliable. In this experiment, we manually annotate the labels of a small set of LLM-generated responses and examine the accuracy of labels generated from ROUGE-L scores. We also use these manual labels to evaluate top-performing algorithms.
To label the data manually, a human labeler checks the true answer of a question and labels the correctness of a response generated by the LLM. Then, automatic labels are compared to manual labels to get the accuracy value. Limited by our resources, we label 600 questions for each of the three datasets, TriviaQA, CoQA, and HotpotQA. Table  4 provides a breakdown of the accuracy across different datasets. The results indicate that labels from ROUGE-L scores are fairly accurate: the lowest accuracy is 0.89. The results are also consistent with previous studies (Kuhn et al., 2022).
Table 4: The accuracy of labels generated from ROUGE-L scores. The high accuracy indicates the reliability of automatic labels, as well as our evaluation above.
TriviaQA CoQA HotpotQA
Accuracy of automatic labels 0.96 0.92 0.89
Table 5: Evaluation with manual labels (Llama3)
Dataset Method AUROC ECE Brier
TriviaQA GS+Platt 0.79 0.050 0.15
Apricot+Platt 0.77 0.059 0.13
Ours 0.81 0.037 0.12
CoQA GS+Platt 0.74 0.056 0.12
Apricot+Platt 0.72 0.062 0.11
Ours 0.76 0.023 0.09
HotpotQA GS+Platt 0.76 0.088 0.19
Apricot+Platt 0.78 0.052 0.18
Ours 0.81 0.030 0.15
Table 6: Evaluation with manual labels (Vicuna)
Dataset Method AUROC ECE Brier
TriviaQA GS+Platt 0.78 0.053 0.18
Apricot+Platt 0.77 0.052 0.17
Ours 0.80 0.032 0.16
CoQA GS+Platt 0.69 0.049 0.19
Apricot+Platt 0.72 0.057 0.15
Ours 0.77 0.032 0.12
HotpotQA GS+Platt 0.68 0.065 0.13
Apricot+Platt 0.77 0.076 0.14
Ours 0.78 0.041 0.12
We further evaluated our model and two strong baselines (GraphSpectral+Platt and Apricot+Platt) with manual labels. As shown in Table 5 and Table 6, our model still outperforms the baselines, which is consistent with our previous evaluation results with automatic labels.
@@ 4.5 Sensitivity analysis
Number of training samples We conducted experiments to examine the relationship between performance and the amount of training data. Specifically, we tested our model performance on the Llama3 TriviaQA dataset and varied the training size from 100 to 4000. The results are displayed in Table  7. We observed that the model’s performance does not drop significantly with the reduced training data. These experimental results indicate that the model performs well with limited data availability, demonstrating its applicability in real-world scenarios where only smaller datasets are available. We also tested the baseline performance, the results are shown in Appendix E.
Hyperparameter sensitivity We conduct the sensitivity analysis of our model’s calibration error performance concerning two key configurations: the number of sampled answers used to construct the graph and the number of Graph Convolutional Network (GCN) layers in the GNN model. The results are displayed in Fig. 3. The experiments are conducted using the Llama3 model on the TriviaQA dataset. For Fig. 3 (a) experiments, we varied the number of sampled answers from 10 to 50 while keeping other configurations and hyperparameters fixed, as described in the experimental setup. We observe that increasing the number of sampled answers slightly improves performance, which then stabilizes. In Fig. 3(b), the sensitivity to the number of GCN layers indicates that our model remains stable with 1 to 4 layers, with the best performance observed at 3 layers.
@@
Tr. size ECE AUROC Brier
100 0.082 0.812 0.201
300 0.062 0.836 0.187
500 0.059 0.842 0.181
1000 0.047 0.853 0.177
4000 0.029 0.860 0.14
@@ 5 Conclusion and Future Work
In summary, in this work, we proposed an effective strategy of calibrating the confidence an LLM’s responses by learning an auxiliary GNN model on the self-consistency pattern among responses to the same questions. Experiments demonstrate that the proposed approach improves confidence calibration significantly across several datasets compared to baseline methods. Our calibration model enhances the reliability of LLMs by evaluating response accuracy, enabling them to abstain from uncertain queries and empowering users to determine trust levels, thereby promoting responsible deployment in society. However, there are instances where an LLM might be highly confident in an incorrect semantic response, resulting in a consistency graph similar to that of a correct answer. In such cases, our calibration model may not provide an accurate confidence estimation. Without a straightforward solution to this problem, we leave the study of this issue to the future.
Broader Impacts:  Our work’s advancement in calibrating the confidence levels of LLM responses carries important social implications: by accurately estimating the likelihood that a response is correct, an LLM can choose to abstain from answering difficult questions or signal its level of confidence to help users make more informed decisions. We believe this work will enhance the trustworthiness of LLMs and contribute positively to their responsible use in society.
Acknowledgment
We thank all reviewers and the action editor for their insightful feedback. Li-Liping Liu’s work was supported by NSF Award 2239869. We would like to thank Xu Han and Jacob Boerma for their help in labeling the data used in this work.
Appendix A Hyperparameters and Model configurations
@@
[…]
@@ Appendix B Computational cost
Appendix C Additional Cases
To better understand our method intuitively, we have collected a few examples to show the difference between our algorithm and APRICOT. To summarize our observation here:
@@ Appendix C Additional Cases
Self-evaluation: 80
@@ Appendix E Additional results
Figure 5: Reliability diagrams for different methods using 10 bins each for TriviaQA from Llama3 model responses. The color and the percentage number within each bar indicate the ratio of responses contained in each bin. Larger values are represented by colors closer to purple.
Figure 6: Reliability diagrams for different methods using 10 bins each for CoQA from Llama3 model responses. The color and the percentage number within each bar indicate the ratio of responses contained in each bin. Larger values are represented by colors closer to purple.
@@
Appendix F Prompting strategy
Here, we showed the prompts to generate the rephrasing questions.
Prompts for rephrasing questions You are a helpful assistant. I have a question that I would like to see it rephrased in multiple ways. Please take the original question and generate several rephrased versions while maintaining the same meaning, and the question can only have one direct answer. Here is the original question: …. Please provide four distinct rephrases of the question.
The prompts for labeling:
Prompts for labeling You will be provided with a question, a reference answer, and a student’s answer. Please evaluate the student’s answer based on the reference answer and provide your score for the student’s answer in the format: “Score: ”. Assign a score of 0 for incorrect and 1 for correct. For example, “Score: 0” or “Score: 1”. Do not include any additional information. Question: {…} Student answer: {…} Reference answer: {…} Now, please enter your score. Score:
Appendix G Sensitivity and ablations
G.1 Sensitivity to accuracy metric
In this section, we evaluate the sensitivity of the threshold of accuracy metric of our models. From the results, it show that our method is relative insensitive of the threshold.
Table 9: The results of using different threshold
Threshold AUROC ECE
Apricot GraphSpectral Ours Apricot GraphSpectral Ours
0.3 0.72 0.79 0.82 0.074 0.076 0.023
0.5 0.70 0.76 0.80 0.091 0.083 0.031
G.2 Use the sentence embedding feature as the node feature
We also provide the comparison with using sentence embedding feature as the node feature. We tested this method on TriviaQA. We got Brier scores 0.21, AUROC 0.75, and ECE values 0.11. The results indicate that GNN with sentence embedding as the node feature can produce worse results than our proposed approach. We see clear overfitting issues when GNN uses semantic features: the validation quickly shoots up after the initial dip. We conclude that GNN using semantic features could not generalize to test data.
G.3 Use the Rouge similarity as the weight of the similarity graph
We provide the results of using Rouge-L as the weight of the similarity graph as shown in Table 10. For the dataset with long-form answers (e.g., TruthfulQA), the performance is much worse than using the clusterID feature,
Table 10: The reuslts of using Rouge similarity graph
ECE AUROC Brier
TriviaQA 0.022 $\pm$ 0.006 0.836 $\pm$ 0.000 0.122 $\pm$ 0.000
CoQA 0.021 $\pm$ 0.007 0.795 $\pm$ 0.001 0.110 $\pm$ 0.000
TruthfulQA 0.035 $\pm$ 0.005 0.632 $\pm$ 0.014 0.221 $\pm$ 0.002
From the results, we conclude Rouge-L is sensitive to the length of the responses.